# MUSTAFAR: Promoting Unstructured Sparsity for KV Cache Pruning in LLM Inference

### v1

@@ 4 Evaluation
Methodology: We evaluate Mustafar on two aspects: Accuracy and Efficiency. For accuracy evaluation, we use tasks from LongBench [3] to test the accuracy retention of per-token magnitude-based pruning of KV cache. We evaluate on three models: Llama-2-7B [34], Llama-3-8B-Instruct [11], and Mistral-7B-Instruct-v0.2 [17]. We also explore the impact of Mustafar when jointly used with orthogonal compression techniques, KV cache quantization KIVI [27] and token-wise eviction H2O [44]. For efficiency evaluation, we evaluate the impact on KV cache compression and computational latency with Llama-2-7B and Llama-3-8B-Instruct. Efficiency evaluation is tested on NVIDIA RTX 6000ADA GPU and measured with NVIDIA Nsight Profiling Tool. Additionally, we provide a accuracy comparison of Mustafar’s unstructured sparisty with 2:4 semi-structured sparsity in Appendix B.
@@ 4.1 LongBench Results
Table 4 shows the extended LongBench evaluation of Mustafar per-token magnitude-based pruning with comparison to dense model and ThinK [38]. Under the same Key cache sparsity, unstructured nature of Mustafar constantly achieves higher accuracy than structured sparsity on ThinK [38] across all tasks. A key advantage of unstructured pruning is its ability to effectively prune the Value cache with minimal accuracy degradation—something that structured pruning has struggled to achieve. Even under high sparsity 70% for both the Key and Value caches, unstructured pruning(yellow) consistently outperforms ThinK’s Key-only 50% structured pruning(pink) on LLaMA-3 8B and Mistral 7B, and achieves comparable accuracy on LLaMA-2-7B.
@@
KV Sparsity NtrvQA Qasper MF-en HotpotQA 2WikiMQA Musique GovReport QMSum MultiNews TREC TrivialQA SAMSum PCount PRe Lec RBP Avg.
@@ 4.2.1 Joint Application with Token Eviction
H2O [44] retains a fixed budget of recent tokens and critical heavy-hitter tokens. Applying Mustafar to H2O, we retain the same scheme of pruning the KV cache of tokens that exit the local dense window. We configure 10% of KV cache budget each to recent tokens and heavy-hitter tokens. Jointly applied, all heavy-hitter tokens and a part of recent tokens is kept as pruned and compressed. We validate the efficacy of Mustafar’s accuracy retention when jointly applied with token eviction, as we see that 50% sparsity in both KV cache retains the dense accuracy with some degradation when pruned to 70% sparsity.
@@ 4.2.2 Joint Application with Quantization
KIVI [27] applies a per-channel quantization of Key cache and per-token quantization of Value cache. Following findings of Harma et al. [13], we first prune each token’s KV cache before quantization is performed. However, we note that current Mustafar Sparse Attention Kernel does not support low-bit precision. Therefore, the accuracy measurement was performed on a sparse quantized KV cache. Table 6 shows the performance of Mustafar and KIVI applied together. Similar to joint application with H2O, we see that model accuracy is retained across the tasks for 50% on Key cache, Value cache, as well as both Key and Value cache. We observe a decrease in accuracy as 70% pruning, with Summarization task seeing the most significant drop. However, other tasks, such as Single-Document QA maintain the same performance as naive 16-bit model, suggesting the potential for applying varying degrees of compression tailored to specific tasks.
@@ 4.3 Efficiency Evaluation
(b) Compression rate-accuracy comparison of Mustafar and ThinK.
Figure 6: Efficiency evaluation of Mustafar. In (b), compression rate refers to percentage of compressed size compared to dense KV cache.
A crucial aspect of Mustafar is to ensure that the exploitation of sparsity for compressing the KV cache does not deter the inference latency. Mustafar compensates the overhead of runtime pruning and compression by achieving speedup in the memory-bound SpMV. Figure 6(a) compares the normalized latency of dense batched MV of cuBLAS with the components of Mustafar Attention Kernel (Figure 5(a)): batched SpMV, dense batched MV of local window, runtime pruning, and compression, for input sequence length 2048 for Llama-2 and 4096 for Llama-3 and generation length 1024. In the multi-head attention of Llama-2-7B, pruning introduces 1.84%, compression introduces 6.25%, and MV of local window introduces 0.62% of the cuBLAS execution time in dense inference. In both 50% and 70%, the speedup gained from SpMV kernel more than compensates for the introduced overheads. In 50% sparsity, SpMV takes 81.07% of cuBLAS execution time and for 70% sparsity, SpMV takes 61.87% of cuBLAS execution time. In Grouped-Query Attention of Llama-3-8B, where there is reduced set of KV cache, compression and pruning overhead reduce down to 1.47% and 0.47% of cuBLAS execution time respectively.
Figure 6(b) compares the KV cache compression rate (of size in memory compared to dense KV cache) of Mustafar and ThinK along with the LongBench average score achieved with Llama-2-7B and Llama-3-8B-Instruct. In this plot, the red arrow points to the optimal direction, where a model achieves higher LongBench score while achieving high compression of the KV cache. For ThinK [38] which prune only Key cache, 50% sparsity leads to 75% compression rate to dense KV cache, and 70% Key cache sparsity leads to 65% compression rate. In the case of Mustafar where both Key and Value Cache can be pruned, KV cache 50% sparsity leads to 65% compression rate. The reason behind 15% additional memory footprint is due to the tile offset overhead as shown in Figure 5(b) and the multiples-of-8 padding enforced to coalesce memory access in GPU. KV cache 70% sparsity leads to 45% compression rate, 50% sparsity to either Key or Value cache leads to 83% compression rate, and single-cache 70% sparsity leads to 72.5% compression rate. Overall, we see that Mustafar is able to achieve better accuracy given the compression rate, with the compression rate-accuracy curve closer to the optimal direction than ThinK.
@@
Figure 7 shows the throughput comparison to inference with dense models. For Llama-2 7B, we used input sequence length of 2048 and generated 2048 tokens. For Llama-3 8B, we use input sequence length of 4096 and generated 4096 tokens. Overall, we see that Mustafar is able to achieve higher throughput as well as support larger batch size owing to the reduced memory footprint of KV cache. In Llama-3, we see that enabling batch size of 8 leads to 2.23$\times$ tokens/sec throughput compared the dense inference of batch size 6. Even within the same batch size, we see an increased throughput upto 1.89$\times$. This is due to the pruning and compression overhead amortized by the speedup of Mustafar sparse attention kernel, leading to faster inference latency. However in batch size 1, we see that throughput is lower than dense inference. This is due to the underutilization GPU in Mustafar sparse attention kernel with small batch size, where the number of threadblocks is smaller than the number of SMs.
@@ 5 Related Work
KV cache compression Alongside aforemention works in KV cache pruning [38, 28], quantization [27, 14, 41, 45], token-wise eviction [44, 26, 22, 7, 1, 10], and low-rank approximation [40, 4, 31, 43, 23], KV cache offloading [21, 25, 12] evicts KV cache to CPU memory and speculatively prefetchs critical tokens’ KV cache. Layer-centric compression [42, 24] applies different level of compression to different layers, adhering to layer-wise importance. Head-level compression [9, 33] applies different level of compression to each heads, from the observation that not all heads contribute equally. Phase-specific compression [36] applies different strategy for prefill and decode phase, with information retention prioritized in prefill and heavy hitter selection applied on decode phase.
System/Kernel for Attention While Mustafar Attention Kernel focuses on operating directly on the bitmap-compressed sparse KV cache, there exists various contributions from the system and kernel-level to optimize for attention. PagedAttention [20] introduces an paging-inspired attention algorithm that partitions KV cache into memory blocks to reduce memory fragmentation and efficient sharing across sequences. FlashDecoding[15] introduces double-buffering do accelerate memory-bound GeMM of decode phase. FlashInfer [39] unifies KV cache format using a block-sparse representation for an efficient management of KV cache that leads to increased throughput. Loki [30] uses a sparse attention method that leverages the low-dimensionality of key vectors to perform an approximate attention in a reduced PCA space.
@@ 6 Conclusion and Limitations
In this work, we demonstrate that unstructured sparsity presents a powerful and novel solution for KV cache pruning. By removing constraints on the pruning pattern, we show that per-token magnitude-based pruning achieves high sparsity while maintaining model accuracy. To unlock the practical benefits of unstructured sparsity, we introduce a bitmap-based sparse format and a custom attention kernel that directly operates on compressed KV cache. Together, our pruning strategy, sparse format, and custom kernel form an end-to-end system that substantially reduces KV cache memory usage and improves throughput, making it possible to support longer contexts and more efficient inference. We believe this work establishes a foundation for future efforts to integrate unstructured sparsity into practical LLM deployment pipelines and opens new directions for memory-efficient LLM inference at scale. Despite our method’s ability to map arbitrary sparsity degree, we constrain ourselves to study the impact of two sparsity degrees: 50% and 70%. While we explore higher sparsity uniformly applied to the entire KV cache in Appendix A.3, a future work involves applying deriving the optimial target sparsity to a smaller granularity, i.e. heads and layers.
@@ A.1 Section 2 Methodology Applied to LLaMA-2 7B
ThinK (Structured) Unstructured Output-aware Unstructured Magnitude ThinK (Structured) Unstructured Output-aware Unstructured Magnitude
@@
ThinK (Structured) Magnitude (Per-channel) Output-aware (Per-channel) Magnitude (Per-token) ThinK (Structured) Magnitude (Per-channel) Output-aware (Per-channel) Magnitude (Per-token)
@@ A.2 Scaling to Larger Model
KV Sparsity NtrvQA Qasper MF-en HotpotQA 2WikiMQA Musique GovReport QMSum MultiNews TREC TrivialQA SAMSum PCount PRe Lec RBP Avg.
@@
A.3 Higher Sparsity
@@
KV Sparsity NtrvQA Qasper MF-en HotpotQA 2WikiMQA Musique GovReport QMSum MultiNews TREC TrivialQA SAMSum PCount PRe Lec RBP Avg.
@@ Appendix B Comparison with semi-structured sparsity
Between the structured pruning of rows and columns, and unstructured pruning of element, lies the 2:4 semi-structured sparsity where 2 out of 4 consecutive elements are non-zero, enforcing a global 50% sparsity. Supported by NVIDIA Sparse Tensor Cores, 2:4 semi-structured sparsity also pursue the same objectives of Mustafar bitmap-based sparse format (Figure 5(b)), maximal compression and fast computation. In Table 12, we apply 2:4 semi-structured pruning to the per-token magnitude-based scheme. Comparing semi-structured sparsity to Key, Value, and both Key and Value cache to unstructured sparsity of Mustafar, we see that unstructured sparsity constantly outperforms semi-structured pattern of the same sparsity. This emphasizes the impact of fine-grained unstructured sparsity of element-wise pruning in model accuracy retention.
@@
KV Sparsity NtrvQA Qasper MF-en HotpotQA 2WikiMQA Musique GovReport QMSum MultiNews TREC TrivialQA SAMSum PCount PRe Lec RBP Avg.
@@
Appendix C Kernel Details
As a supplement to Section 3, we offer more detail onto the Mustafar Sparse Attention Kernel, which accelerates memory-bound SpMV.
C.0.1 Load-as-Compressed, Compute-as-Dense Pipeline
@@
Crucial insight of accelerating SpMV involves reducing the data movement between the GPU global memory and the local memory of each GPU Streaming Multiprocessor. First proposed by FlashLLM [37], load-as-compressed, compute-as-dense pipeline as shown in Figure 8 involves sending each matrix tile in the corresponding compressed form to the SM registers (’gmem2reg’ in the figure), decompressing the compressed tile into the dense from to the shared memory (’extract’), then initializing computation on the next pipeline stage (’smem2tc’). Computation is mapped to tensor core to utilize the high fp16 compute throughput. To map MV, unused N dimensions are padded to zero for computation. Non-zero thread-tile of $1\times 64$ in Figure 5(b) represents the granularity of non-zeros that a warp thread decompresses at a pipeline stage. Each warp thread decompresses 2 thread-tile per stage using the corresponding bitmap to determine the correct position of each non-zero. Effectively, each warp operates on a $64\times 64$ sized matrix tile at a time.
C.0.2 KV cache Management

### v2

@@ 4 Evaluation
Methodology: We evaluate Mustafar on two aspects: Accuracy and Efficiency. For accuracy evaluation, we use tasks from LongBench [3] to test the accuracy retention of per-token magnitude-based pruning of KV cache. We evaluate on three models: Llama-2-7B [40], Llama-3-8B-Instruct [12], and Mistral-7B-Instruct-v0.2 [19]. We also explore the impact of Mustafar when jointly used with orthogonal compression techniques, KV cache quantization KIVI [30] and token-wise eviction H2O [51]. For efficiency evaluation, we evaluate the impact on KV cache compression and computational latency with Llama-2-7B and Llama-3-8B-Instruct. Efficiency evaluation is tested on NVIDIA RTX 6000ADA GPU and measured with NVIDIA Nsight Profiling Tool. Additionally, we provide accuracy evaluation on RULER [17] benchmark in Appendix A.3, accuracy comparison of Mustafar’s unstructured sparsity with 2:4 semi-structured sparsity in Appendix B, and additional kernel throughput evaluation in Appendix C.3.
@@ 4.1 LongBench Results
Table 4 shows the extended LongBench evaluation of Mustafar per-token magnitude-based pruning with comparison to dense model and ThinK [44]. Under the same Key cache sparsity, unstructured nature of Mustafar constantly achieves higher accuracy than structured sparsity on ThinK [44] across all tasks. A key advantage of unstructured pruning is its ability to effectively prune the Value cache with minimal accuracy degradation, which structured pruning has struggled to achieve. Even under high sparsity 70% for both the Key and Value caches, unstructured pruning (yellow) consistently outperforms ThinK’s Key-only 50% structured pruning (pink) on LLaMA-3 8B and Mistral 7B, and achieves comparable accuracy on LLaMA-2-7B.
@@
KV
Sparsity
NtrvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TrivialQA
SAMSum
PCount
PRe
Lec
RBP
Avg.
@@ 4.2.1 Joint Application with Token Eviction
H2O [51] retains a fixed budget of recent tokens and critical heavy-hitter tokens. Applying Mustafar to H2O, we retain the same scheme of pruning the KV cache of tokens that exit the local dense window. We configure 10% of KV cache budget each to recent tokens and heavy-hitter tokens. Jointly applied, all heavy-hitter tokens and a part of recent tokens is kept as pruned and compressed. In Table 5, we validate the efficacy of Mustafar’s accuracy retention when jointly applied with token eviction, as we see that 50% sparsity in both KV cache retains the dense accuracy with some degradation when pruned to 70% sparsity.
@@ 4.2.2 Joint Application with Quantization
KIVI [30] applies a per-channel quantization of Key cache and per-token quantization of Value cache. Following findings of Harma et al. [14], we first prune each token’s KV cache before quantization is performed. However, we note that current Mustafar sparse attention kernel implementation does not support low-bit precision. Therefore, the accuracy measurement was performed on a sparse quantized KV cache. Table 6 shows the performance of Mustafar and KIVI applied together. Similar to joint application with H2O, we see that model accuracy is retained across the tasks for 50% on Key cache, Value cache, as well as both Key and Value caches. We observe a decrease in accuracy at 70% pruning, with Summarization task seeing the most significant drop. However, other tasks, such as Single-Document QA maintain the same performance as naive 16-bit model, suggesting the potential for applying varying degrees of compression tailored to specific tasks.
@@ 4.3 Efficiency Evaluation
(b) Compression ratio-accuracy comparison of Mustafar and ThinK.
Figure 6: Efficiency evaluation of Mustafar. In (b), compression ratio refers to percentage of compressed size compared to dense KV cache.
A crucial aspect of Mustafar is to ensure that the exploitation of sparsity for compressing the KV cache does not deter the inference latency. Mustafar compensates the overhead of runtime pruning and compression by achieving speedup in the memory-bound SpMV. Figure 6(a) compares the normalized latency of dense batched MV of cuBLAS with the components of Mustafar sparse attention kernel (Figure 5(b)): batched SpMV, dense batched MV of local window, runtime pruning, and compression, for input sequence length 2048 for Llama-2 and 4096 for Llama-3 and generation length 1024. In the multi-head attention of Llama-2-7B, pruning introduces 1.84%, compression introduces 6.25%, and MV of local window introduces 0.62% of the cuBLAS execution time in dense inference. In both 50% and 70%, the speedup gained from SpMV kernel more than compensates for the introduced overheads. In 50% sparsity, SpMV takes 81.07% of cuBLAS execution time and for 70% sparsity, SpMV takes 61.87% of cuBLAS execution time. In Grouped-Query Attention of Llama-3-8B, where there is reduced set of KV cache, compression and pruning overhead reduce down to 1.47% and 0.47% of cuBLAS execution time respectively.
Figure 6(b) compares the KV cache compression ratio (% of size in memory compared to dense KV cache) of Mustafar and ThinK along with the LongBench average score achieved with Llama-2-7B and Llama-3-8B-Instruct. In this plot, the red arrow points to the optimal direction, where a model achieves higher LongBench score while achieving high compression of the KV cache. For ThinK [44] which prune only Key cache, 50% sparsity leads to 75% compression ratio to dense KV cache, and 70% Key cache sparsity leads to 65% compression ratio. In the case of Mustafar where both Key and Value Cache can be pruned, KV cache 50% sparsity leads to 65% compression ratio. The reason behind 15% additional memory footprint is due to the tile offset overhead as shown in Figure 5(a) and the multiples-of-8 padding enforced to coalesce memory access in GPU. KV cache 70% sparsity leads to 45% compression ratio, 50% sparsity to either Key or Value cache leads to 83% compression ratio, and single-cache 70% sparsity leads to 72.5% compression ratio. Overall, we see that Mustafar is able to achieve better accuracy given the compression ratio, with the compression ratio-accuracy curve closer to the optimal direction than ThinK.
@@
[…]
@@ 5 Related Work
KV cache compression Alongside aforementioned work in KV cache pruning [44, 31], quantization [30, 15, 48, 52], token-wise eviction [51, 29, 25, 8, 1, 11], and low-rank approximation [47, 4, 37, 50, 26], KV cache offloading [24, 28, 13, 5] evicts KV cache to CPU memory and speculatively prefetchs critical tokens’ KV cache. Layer-centric compression [49, 27] applies different level of compression to different layers, adhering to layer-wise importance. Head-level compression [10, 39] applies different level of compression to each heads, from the observation that not all heads contribute equally. Phase-specific compression [42] applies different strategy for prefill and decode phase, with information retention prioritized in prefill and heavy hitter selection applied on decode phase.
System/Kernel for Attention While Mustafar attention kernel focuses on operating directly on the bitmap-compressed sparse KV cache, there exists various contributions from the system and kernel-levels to optimize for attention. PagedAttention [23] introduces an paging-inspired attention algorithm that partitions KV cache into memory blocks to reduce memory fragmentation and efficient sharing across sequences. FlashDecoding[16] introduces double-buffering to accelerate memory-bound GeMM of decode phase. FlashInfer [45] unifies KV cache format using a block-sparse representation for an efficient management of KV cache that leads to increased throughput. Loki [36] uses a sparse attention method that leverages the low-dimensionality of key vectors to perform an approximate attention in a reduced PCA space.
@@ 6 Conclusion and Limitations
In this work, we demonstrate that unstructured sparsity presents a powerful and novel solution for KV cache pruning. By removing constraints on the pruning pattern, we show that per-token magnitude-based pruning achieves high sparsity while maintaining model accuracy. To unlock the practical benefits of unstructured sparsity, we introduce a bitmap-based sparse format and a custom attention kernel that directly operates on compressed KV cache. Together, our pruning strategy, sparse format, and custom kernel form an end-to-end system that substantially reduces KV cache memory usage and improves throughput, making it possible to support longer contexts and more efficient inference. Mustafar establishes a foundation for future efforts to integrate unstructured sparsity into practical LLM deployment pipelines and opens new directions for memory-efficient LLM inference at scale. In future work, we plan to explore the joint effect of leveraging KV sparsity of Mustafar with sparsity in weights derived by works such as output-aware weight pruning [38], pruning with low-rank adapters for accuracy retention [34, 33], and activation-aware calibration and efficiency enhancement [32, 46]. Additionally, this paper focuses on showing that unstructure sparsity can prune both Key and Value caches to a higher sparsity with better accuracy than structured sparsity, leaving our method’s ability to map arbitrary sparsity degree untouched. While we explore higher sparsity uniformly applied to the entire KV cache in Appendix A.4, a future work involves deriving the optimal target sparsity to a smaller granularity (e.g. per-head or per-layer) to maximize sparsity and accuracy retention.
Acknowledgments and Disclosure of Funding
We gratefully acknowledge the support of National Science Foundation (NSF) under program PPoSS, Award Number 2316177.
@@ A.1 Section 2 Methodology Applied to LLaMA-2 7B
ThinK
Unstructured
Output-aware
Unstructured
Magnitude
ThinK
Unstructured
Output-aware
Unstructured
Magnitude
@@
ThinK
Magnitude
(Per-channel)
Output-aware
(Per-channel)
Magnitude
(Per-token)
ThinK
Magnitude
(Per-channel)
Output-aware
(Per-channel)
Magnitude
(Per-token)
@@ A.2 Scaling to Larger Model
KV
Sparsity
NtrvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TrivialQA
SAMSum
PCount
PRe
Lec
RBP
Avg.
@@
A.3 Evaluation on RULER
For a more diverse evaluation, we evaluate Llama-3.1-8B-Instruct on RULER [17] benchmark for context length of 65,536 tokens.
Table 11: Accuracy comparison on RULER benchmark
Sparsity
Method
Needle-Single1
Needle-Single2
Needle-MultiKey1
Needle-MultiKey2
Needle-MultiQuery
Needle-MultiValue
QA-1
QA-2
Variable Tracking
Freq. Words Extract.
Llama-3.1-8B-Instruct
Dense 1.000 1.000 0.990 0.979 0.990 0.979 0.844 0.594 0.973 0.851
Key 50% ThinK 1.000 1.000 0.990 0.979 0.995 0.969 0.833 0.594 0.919 0.854
Mustafar 1.000 1.000 0.990 0.979 0.995 0.996 0.833 0.573 0.971 0.813
Key 70% ThinK 0.448 0.490 0.229 0.188 0.646 0.487 0.615 0.510 0.208 0.427
Mustafar 1.000 1.000 0.990 0.969 0.992 0.903 0.833 0.594 0.966 0.823
Value 50% ThinK 1.000 1.000 0.990 0.969 0.914 0.958 0.823 0.573 0.910 0.792
Mustafar 1.000 1.000 0.979 0.995 0.995 0.971 0.833 0.604 0.983 0.830
Value 70% ThinK 0.948 0.927 0.948 0.510 0.698 0.688 0.646 0.500 0.558 0.677
Mustafar 1.000 1.000 1.000 0.979 0.992 0.969 0.833 0.594 0.985 0.826
Key&Value 50% ThinK 0.958 1.000 0.948 0.854 0.828 0.956 0.740 0.531 0.742 0.823
Mustafar 1.000 1.000 0.990 0.979 0.997 0.997 0.833 0.573 0.862 0.809
Key&Value 70% ThinK 0.000 0.073 0.000 0.000 0.000 0.000 0.219 0.250 0.000 0.035
Mustafar 1.000 1.000 0.990 0.969 0.995 0.914 0.833 0.583 0.869 0.799
As shown in Table 11, even in the challenging Needle-in-a-Haystack scenarios with multiple keys and queries, Mustafar maintains accuracy comparable to the dense model. It also outperforms the structured pruning baseline ThinK, with particularly notable gains at 70% joint Key-Value sparsity. While structured pruning does perform well in isolated cases, such as the Needle-Single tasks for 70% Value sparsity, it exhibits significant accuracy drops in other tasks. In contrast, Mustafar’s unstructured sparsity consistently preserves accuracy across all tasks. This contrast highlights the versatility of unstructured sparsity in adapting to diverse task requirements.
A.4 Higher Sparsity
@@
KV
Sparsity
NtrvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TrivialQA
SAMSum
PCount
PRe
Lec
RBP
Avg.
@@ Appendix B Comparison with Semi-structured Sparsity
Between the structured pruning of rows and columns, and unstructured pruning of element, lies the 2:4 semi-structured sparsity where 2 out of 4 consecutive elements are non-zero, enforcing a global 50% sparsity. Supported by NVIDIA Sparse Tensor Cores, 2:4 semi-structured sparsity also pursue the same objectives of Mustafar bitmap-based sparse format (Figure 5(a)), maximal compression and fast computation. In Table 13, we apply 2:4 semi-structured pruning to the per-token magnitude-based scheme. Comparing semi-structured sparsity to Key, Value, and both Key and Value cache to unstructured sparsity of Mustafar, we see that unstructured sparsity constantly outperforms semi-structured pattern of the same sparsity. This emphasizes the impact of fine-grained unstructured sparsity of element-wise pruning in model accuracy retention.
@@
KV
Sparsity
NtrvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TrivialQA
SAMSum
PCount
PRe
Lec
RBP
Avg.
@@
Appendix C Sparse Attention Kernel Details
As a supplement to Section 3, we offer more detail onto the Mustafar sparse attention kernel, which accelerates memory-bound batch SpMV.
C.1 Load-as-Compressed, Compute-as-Dense Pipeline
@@
Crucial insight of accelerating SpMV involves reducing the data movement between the GPU global memory and the local memory of each GPU Streaming Multiprocessor. First proposed by FlashLLM [43], load-as-compressed, compute-as-dense pipeline as shown in Figure 8 involves sending each matrix tile in the corresponding compressed form to the SM registers (’gmem2reg’ in the figure), decompressing the compressed tile into the dense from to the shared memory (’extract’), then initializing computation on the next pipeline stage (’smem2tc’). Computation is mapped to tensor core to utilize the high fp16 compute throughput. To map MV, unused N dimensions are padded to zero for computation. Non-zero thread-tile of $1\times 64$ in Figure 5(a) represents the granularity of non-zeros that a warp thread decompresses at a pipeline stage. Each warp thread decompresses 2 thread-tile per stage using the corresponding bitmap to determine the correct position of each non-zero. Effectively, each warp operates on a $64\times 64$ sized matrix tile at a time.
C.2 KV Cache Management